OpenStudy (anonymous):

Write an equation that meets the given conditions? A line that passes through the points (2,7) and (6,15) A line that passes through (2,1) and (6,9)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 15}})\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\qquad \textit{plug in your values and solve for "y"}\\ \qquad \uparrow \\ \textit{point-slope form}\)

OpenStudy (anonymous):

alternatively, y = mx + b substitute your known x, y values to find m and b